vidyo link

Action Items:


  • very few persons

Framework & production

  • Federica's position to be checked for next year


Dibosons, VZ resonance (V=W,Z)

  • on CWR comments

B physics

  • (Alberto) studied number of events seen in the trigger. 20% increase for L unit in both triggers. To be checked the superposition of the two (50% max). In contact with TMVA experts (see last week twiki).
  • working with BPH muon POG to set as official CMS tool the new MVA identification developed for BsMuMu.

Dark Matter

  • dark Higgs is in progress. Still working on ntuples code and technical support needed from M.Cremonesi (Thanksgiving last week).
  • some examples run on Padova clusters (spark)

hh4b (non resonant)

  • preApproval talk last week. Several comments we have to address but no evident show-stopper.
  • plan to update documentation within this week.

Same sign dileptons

  • Ugo showed results last week.

Multiple leptons


Flavor tagging

  • (Pablo) bugs fixing on-going on the pull request

*(Martino) pixelDOC shifts done at p5. several tests on dc-dc converters during interfill (even at night). issue still not fully solved. Plan is to extract FPIX+ within Dec.23rd.

*(Paul) update on 2016 lumi mearurements should come soon. Measure of the 2017 lumi within end of the year. Uncertainty about the 2016 one

*(Siew) start working for GIF++. Need to check amount of EPR.




-- Main.Franco Simonetto 04-10-2017

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2017-11-27 - MartinoDallOsso
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback